a college student invested 5000.00 total in 2 seperate accounts. He invested half into a mutual account and the rest into a stock that yielded four times as much interest. After one year he earned a total of 250.00 in interest. What was the interest rates of the mutual fund and stock account?
Answer by rfer(16322)

total interest 250, 4 times more in one than the other, 250=5/5 or 50=1/5
50/2500=.02=2%
200/2500=.08=8%